Circular No. 7099

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990123 J. Heise, Space Research Organization of the Netherlands, Utrecht;

C. DeLibero, M. R. Daniele, and G. Scotti, BeppoSAX Operation Control Center, Telespazio, Rome; D. Ricci and M. Capalbi, BeppoSAX Science Data Center, Rome; L. A. Antonelli, Osservatorio Astronomico, Rome; and E. Costa, Istituto di Astrofisica Spaziale, CNR, Rome, report: “The BeppoSAX Wide Field Camera (WFC) error box of GRB 990123 (IAUC 7095) was observed from Jan. 23.65 UT (6 hr after the γ-ray trigger) until ∼ Jan. 24.75, using the Narrow Field Instruments of BeppoSAX. A bright, previously unknown x-ray source designated as 1SAX J1525.5+4446, was detected by the LECS and MECS units at α = 15^{h}25^{m}31^{s}, δ = +44^{o}46′.3 (equinox 2000.0), with an error-circle radius of 50′′. The new source is within the WFC error circle and at a distance of 22′′ from the optical transient reported on IAUC 7094. The intensity of the source at 2–10 keV in the MECS at the start of the observation is 0.14 count/s, corresponding to 1.1 × 10^{−11} erg cm^{−2} s^{−1}. The source fades by an order of magnitude in 24 hr, consistent with a decay in time with a power-law index of −1.35. The average featureless spectrum has a photon power-law index of ∼ 2, with an interstellar absorption column of 1.2 × 10^{21} cm^{−2}, above the galactic absorption of 2.1 × 10^{20} cm^{−2} expected in that direction. We conclude that we have observed the x-ray afterglow of GRB 990123. It is the brightest of all γ-ray-burst x-ray afterglows observed until now.”

A. Maury, Observatoire de la Cote d’Azur, Caussols; and M. Boer and S. Chaty, Centre d’Etude Spatiale des Rayonnements, Toulouse, France, report: “We clearly detected the optical candidate for GRB 990123 and its presumed host galaxy with the Observatoire de la Côte d’Azur 0.9m Schmidt telescope, in a series of 300-s unfiltered frames. The unfiltered magnitudes (mean uncertainty 0.4 mag; USNO A2.0 reference stars) correspond approximately to R magnitudes, given the detector response: Jan. 24.03703 UT, 19.6; 24.05241, 19.6; 24.05723, 20.0; 24.06188, 19.8; 24.06653, 19.7; 24.07119, 19.4; 24.07584, 20.4; 24.08049, 19.9; 24.08515, 20.1; 24.09445, 20.3; 24.09910, 20.2; 24.11306, 19.7; 24.11773, 19.8; 24.12237, 20.2; 24.12702, 20.5; 24.13167, 20.3; 24.13633, 20.5; 24.14098, 20.2; 24.14563, 20.0; 24.15028, 20.1; 24.15494, 20.4; 24.15959, 20.4; 24.16424, 20.0; 24.16890, 20.1; 24.17820, 20.2; 24.18285, 20.1; 24.20107, 20.4. The resulting decay slope, 1.35, is in good agreement with the slope mentioned by Bloom et al. (GCN 208), given the uncertainties.”

1999 January 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green